Systems Staff Meeting Notes 1/12/98

Systems Staff Meeting Notes
January 12, 1998

Present: Rich Bennett, Bill Covey, Destry Fudge, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Library staff should be returning from ALA by Wednesday, January 14.
  2. Student Schedule - a gap still exists during the lunch hour, may ask Jason Monahan to come in during these hours, as he is here until February 12.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Jason has been checking on incoming items that should be here but haven’t been located. Gresh has been looking for parts for Education.
  2. Parts to upgrade work stations - have received some, not all.
  3. MSL power outage - Lee brought everything up OK on Sunday - lost only two Telex terminals. The NT5 stayed down until staff brought it up today. Pat Haskins had a printer problem, which required a recycle of the ASANTE print box.
Liaisons
  1. Taking attendance has increased the number of attendees.
  2. Vast numbers of them have converted their email - storage on the server is picking up 2 to 3 M per day.
Networking
  1. Jason is checking on the switches - probably somewhere between us and CISCO.
  2. CORE remains the same - 2 rings.
  3. Student housing will supply a network connection for every pillow - mandatory. Will have a $1.2 million budget - looking at 3COM & CISCO.
  4. New network work station software not yet received.
  5. New wiring in LAC - in two weeks.
  6. Gresh is working on Education.
  7. Systems wiring on hold.
Software Projects Status
  1. OPS tracking - list of sample functions given to Winston this morning - will want to integrate with Personnel files.
  2. Invoices delivered to Accounting by Winston this morning. He will deliver the server version to them by tomorrow.
Systems Move
  1. On hold for more electric & phone installation.
  2. Sharon points out that the sign on the outside of the back stairwell says, ‘Stairwell #8’, but the sign on the inside says, ‘Stairwell #7’.
  3. Need confirmation on the working of the electromagnetic locks on the outside doors.
  4. New circuits in the server room - with an accessible breaker may be installed - will cost extra.
  5. Closure on the back and south side doors is not working properly.
  6. The caged area hallway should be cleared out sometime this week.
  7. Student lockers need to be re-keyed.
Calendar
  1. Winston out this afternoon.
  2. Destry back from ALA on Tuesday, Rich on Wednesday.
  3. Martin Luther King, Jr. Holiday - Monday, Jan. 19.
  4. Bill out January 29.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu